A payday loan company charges 2.2 percent interest for a two-week period. What would be the annual interest rate from that company? (Assume an even 52 weeks per year. Enter your answer as a percent rounded to 1 decimal place.)

Annual Interest Rate -
Answer #1

Effective annual rate (EAR), is mathematically represented as:

EAR = (1-6)*-1 i = stated annual interest rate n = number of compounding periods

EAR = (1 +0.022 26 - 1

EAR = 1.7609 - 1

EAR = 0.7609

EAR 76.1%